OpenStudy (anonymous):

What is the solution? 1/2x2 – x + 5 = 0

OpenStudy (anonymous):

Factor out the 1/2. Times each side by 1/2.

OpenStudy (anonymous):

You should be able to see it then :)

OpenStudy (anonymous):

A. 1 ± i sqrt 11 B. –1 ± i sqrt 9 C. 1 ± i sqrt 9 D. –1 ± i sqty 11

OpenStudy (anonymous):

Use equation:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] where you have an equation in the form ax^2 + bx + c = 0. Pay close attention to the determinant b^2-4ac to see whether you have real solutions or complex solutions.

OpenStudy (anonymous):

how do i do this

OpenStudy (anonymous):

ax^2 + bx + c is the formula that you use to find the variables needed for the quadratic equation

OpenStudy (anonymous):

a=1/2, b=-1, and c=5, from the equation in the problem and the form it fits of ax^2 + bx + c = 0. Substitute those values of a, b, and c that I identified for you into the equation in my first post.

OpenStudy (anonymous):

How are you coming along with this? Without simplifying, are you able at least to do the substitutions?

OpenStudy (anonymous):

so the answer will be C. 1 ± i sqrt 9 ?

OpenStudy (anonymous):

Yes! You got it! Good job!

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

You're welcome! Nice working with you!